# V2I-Calib: A Novel Calibration Approach for Collaborative Vehicle and Infrastructure LiDAR Systems

### v1

@@ III-D Extrinsics Resolution and Optimization
The aim of this section is to compute the extrinsic parameters from one or more pairs of matched detection boxes and to ensure the accuracy of these parameters. This process confronts two major challenges: first, the correspondence of the eight point pairs within the matched detection boxes is unknown, as this information is affected by the rotation matrix of the extrinsic parameters; second, the calibration of extrinsic parameters presents itself as an inherently ill-posed issue, necessitating an good initial estimate to strike a balance between accuracy and computational expense.
To address these challenges, let =$\mathcal{B}=\{(\mathbf{B}_{i}^{\text{(inf)}},\mathbf{B}_{i}^{\text{(veh)}})\}_{i=1}^{n}$ denote the set of matched detection box pairs, where $\mathbf{B}_{i}^{\text{(inf)}}$ and $\mathbf{B}_{i}^{\text{(veh)}}$ are the detection boxes from the infrastructure and vehicle, respectively. Based on Equation 3, we derive abstract point clouds $\mathbf{P}^{\text{(inf)}}$ and $\mathbf{P}^{\text{(veh)}}$, which encapsulate the vertex data from matched detection boxes. This transformation redefines the initial cross-source point cloud calibration challenge into a standardized point set registration problem.
Figure 4: Extrinsic Resolution Process. This schematic depicts the stages involved in deriving extrinsic parameters from matched detection boxes. The diagram illustrates the challenges of partial correspondence within the two points lists.
The abstract point clouds exhibit a unique partial correspondence relationship. Specifically, for each $i$, the group created by the bijective mapping between the vertices of $\mathbf{B}_{i}^{\text{(inf)}}$ and $\mathbf{B}_{i}^{\text{(veh)}}$ sets up a group-level correspondence. Nonetheless, the internal correspondence among the point pairs within $\mathcal{C}_{i}$ remains unguaranteed.
$\forall k\in\{8i,\dots,8i+7\},\quad\not\exists\,\mathbf{p}^{\text{(inf)}}_{k}\leftrightarrow\mathbf{p}^{\text{(veh)}}_{k}$ (9)
where $\mathbf{p}^{\text{(inf)}}_{k}$ and $\mathbf{p}^{\text{(veh)}}_{k}$ are the points in $\mathbf{P}^{\text{(inf)}}$ and $\mathbf{P}^{\text{(veh)}}$ respectively, originating from the vertices of the $i$-th matched detection box pair in the dataset.
Based on this characteristic, applying a naive point cloud extrinsic parameter calculation method [19] can provide an rough initial value of the extrinsics. Subsequently, based on this initial value, extrinsic parameter optimization is conducted [24][25] to derive the optimal extrinsic parameter values.
@@ IV EXPERIMENT
In this chapter, we will provide a detailed overview of the experimental design, evaluation metrics, dataset characteristics, and specific experimental settings for assessing the performance of our proposed V2I-Calib method.
@@ IV-A Evaluation Metrics
$\theta_{\text{degrees}}=\arccos\left(\frac{\text{tr}(\mathbf{\Delta R})-1}{2}\right)\times\frac{180}{\pi}$ (10)
@@
$\emph{RTE}=||\mathbf{t}_{t}^{-1}-\mathbf{t}_{e}||_{2}$ (12)
@@ IV-B Datasets
In this study, we conducted experimental validation using the DAIR-V2X dataset[26], which encompasses a wealth of data collected from Vehicle-Infrastructure Cooperative Autonomous Driving (VICAD) scenarios, including vehicular and infrastructural LiDAR data along with their 3D box annotations. The specifications of the vehicular and roadside LiDAR systems are presented in Table I.
@@
To enhance the assessment of our method’s effectiveness across varied scenarios, we have instituted a data difficulty categorization. While conventional perception datasets like KITTI have established a difficulty rating based on factors such as the count, dimensions, and levels of occlusion or truncation of objects, similar frameworks for calibration tasks within publicly available datasets are notably absent. Mindful of scenario classifications in extant literature and the absence of established difficulty metrics for calibration, we have methodically classified the DAIR-V2X dataset into distinct levels of difficulty, making the criteria and associated experimental code readily accessible.
Our approach to defining data difficulty draws from the established logic used in perception tasks, factoring in the intricacy of the environment, the spatial relation between vehicles and infrastructure, and the presence of shared objects within the scene. For instance, we utilized categories from the KITTI dataset[27]—such as object count, size, and states of occlusion or truncation—as a basis to evaluate their influence on the calibration process for LiDAR systems between vehicles and infrastructure. It’s important to note that we also accounted for the effect of relative positioning within vehicular-roadside scenarios and the possibility of missing shared objects in counter scenarios, highlighting that a shared zone in point clouds can enable the gathering of required shared objects by broadening the detection range, albeit potentially at the expense of increased detection time.
@@ IV-C Experimental Settings
In this section, a systematic performance evaluation of the proposed method will be conducted through comparative experiments and ablation studies. Contrast experiments are aimed at verifying the superiority of the proposed method by comparing its performance with other methods on datasets of varying difficulty. On the other hand, ablation studies further analyze the effectiveness of individual components/modules. Both experiments were conducted on an experimental platform equipped with an Intel i7-9750H CPU.
@@ IV-C1 Contrast Experiment
The comparative experiments assessed the performance of the method presented in this article against other existing methods on metrics such as Rotation Error (RRE), Translation Error (RTE), and success rate. The existing methods are primarily categorized into traditional approaches represented by VGICP [12] and NDT[25], and perception-based approaches exemplified by VIPS [7] and HPCR-VI [10].
@@
easy group VGICP [12] 3.18 6.43 6.0 0.77
NDT [25] 1.88 6.36 6.4 19.82
HPCR-VI [10] 5.15 0.91 81.1 /
VIPS [7] 0.96 0.84 79.3 0.33
@@
hard group VGICP [12] 3.11 6.91 4.1 0.42
NDT [25] 2.33 3.91 6.0 29.62
HPCR-VI [10] 6.43 1.52 43.3 /
VIPS [7] 2.12 2.05 57.5 0.29
@@
As indicated in Table II, it is apparent that traditional point cloud registration algorithms largely fail on the DAIR-V2X dataset, and the existing perception-based methods exhibit significant errors. In contrast, the method proposed in this article achieves better results in terms of both precision and real-time performance. Additionally, by incorporating detection box confidence in the engineering implementation, our method effectively mitigates the problem of calibration precision degradation under multi-target interference compared to [10]. In terms of immediacy, it meets the calibration time requirement of less than 0.35 seconds for common intersection traffic scenarios as analyzed in [18]. Notably, the running times for the methods in the more challenging group are generally lower than those in the simpler group, which is attributable to the fewer shared targets in complex scenes, reducing the computational load.
This series of experimental validations confirms that the proposed vehicle-to-infrastructure calibration method not only provides highly accurate calibration results across various scenarios but also demonstrates its feasibility in terms of immediacy, offering robust support for the practical application of vehicle-to-infrastructure LiDAR systems.
(a) Ideal Scenario
(b) Challenging Scenario
(c) Complex Scenario
(d) Complex Scenario
(a) RE Comprison
(b) TE Comprison
(c) Time Cost Comprison
Figure 6: Ablation Comparison of Different Affinity Strategies. The violin plots correspond to the left-side metrics of TE, RE, and Time Cost, while the line charts relate to the success rate metric on the right. We compared four strategies for constructing the affinity matrix based on different combinations of affinities: Strategy 1 combines angle affinity with category affinity, Strategy 2 combines core affinity with category affinity, Strategy 3 employs core affinity alone, and Strategy 4 combines length affinity with category affinity. It is evident that Strategy 2, utilized by our method, surpasses the other strategies in terms of both extrinsic parameter accuracy and computational time.
@@ IV-C2 Ablation experiment
In this section, we aim to validate the effectiveness of the core affinity formulation module presented in Section III-B and the extrinsic parameter optimization module discussed in Section III-D through a series of ablation studies.
@@
Comparing V2I-Calib-v1 with V2I-Calib-v3, we observe that the core affinity formulation module proposed in this paper exhibits enhanced robustness, achieving an approximate 20% increase in the success rate compare. Moreover, we evaluated the impact of the extrinsic parameter optimization module on calibration outcomes. A comparative analysis between V2I-Calib-v1 and V2I-Calib-v2, before and after the introduction of the extrinsic optimization module, reveals that the optimization algorithm significantly improves the precision and success rate of the calibration.
To validate the superiority of the affinity metrics proposed in this paper, we conducted comparative analyses with category affinity, angle affinity, and length affinity as delineated in [7]. The results, as depicted in Figure 6, demonstrate that Strategy 2, which is core affinity, achieves a higher success rate and extrinsic parameter accuracy within an acceptable time frame. Notably, when comparing Strategy 2 (core affinity with category affinity) with Strategy 3 (core affinity without category affinity), it is apparent that the inclusion of category affinity significantly reduces the overall computational time and notably enhances performance within the hard group.
Regarding the Strategy 1 angle affinity and Strategy 4 length affinity extended from [7], these strategies were originally designed for scenarios assisted by positional information. However, experimental results show that they are somewhat applicable in scenarios without initial extrinsic parameter values, although their performance significantly decreases in challenging settings. Upon further analysis of the scenarios where these strategies failed, we observed that most of these scenarios involved extrinsic parameters with relatively large absolute values, aligning with the limitations inherent in their initial design rationale.
These ablation studies not only confirm the efficacy of each distinct module within our framework but also underscore their collective role in enhancing the overall functionality of the system.
@@ V CONCLUSIONS
This study introduces a novel calibration method, V2I-Calib, for cooperative vehicle and infrastructure LiDAR systems. Through a series of experiments, our method demonstrates superior calibration accuracy and robustness across datasets of various difficulties, particularly excelling in handling complex scenes while maintaining real-time performance. In essence, the calibration task for cooperative vehicle-infrastructure LiDAR systems is fundamentally a data association task, bearing some similarity to multi-object tracking tasks.
Future research will integrate the concept of temporal association from multi-object detection, merging spatial alignment and temporal synchronization tasks. Furthermore, the integration of 2D bounding box information will shift the research focus towards multi-sensor fusion calibration in cooperative vehicle-infrastructure systems.
As cooperative vehicle-infrastructure technology continues to evolve, we anticipate this framework to adapt to more application scenarios, providing a solid foundation for the safety and reliability of autonomous driving systems.

### v2

@@ III-C Extrinsics Resolution and Optimization
The objective of this section is to compute extrinsic parameters from matched detection box pairs and to ensure their accuracy, addressing two primary challenges: the unknown correspondence of point pairs within the boxes, influenced by the rotation matrix of the extrinsics, and the ill-posed nature of calibration that requires a good initial estimate to balance accuracy with computational cost.
After obtaining $\mathcal{B}^{(c)}$, we derive the feature point clouds $\mathbf{P}^{\text{(inf)}}$ and $\mathbf{P}^{\text{(veh)}}$ using Equation (3), transforming the initial cross-source point cloud calibration challenge into a standardized point set registration problem.
A naive extrinsic parameter calculation method provides an initial value [18], which is optimized to derive the optimal parameters [23, 24].
@@ IV-A Evaluation Metrics
$\emph{RRE}=\arccos\left(\frac{\text{tr}(\mathbf{\Delta R})-1}{2}\right)$ (8)
@@
$\emph{RTE}=||\mathbf{t}_{t}-\mathbf{t}_{e}||_{2}$ (10)
@@ IV-B Datasets
In this study, we conducted experimental validation using the DAIR-V2X dataset[25], which encompasses a wealth of data collected from Vehicle-Infrastructure Cooperative Autonomous Driving (VICAD) scenarios, including vehicular and infrastructural LiDAR data along with their 3D box annotations. The specifications of the vehicular and roadside LiDAR systems are presented in Table I. The outcomes of the experimental scenarios are illustrated in Figure 5.
@@ IV-C Experimental Settings
This section describes a systematic performance evaluation of the proposed method, which is conducted through comparative experiments and ablation studies.All experiments were performed on an experimental platform equipped with an Intel i7-9750H CPU.
@@ IV-C1 Contrast Experiment
The comparative experiments assessed the performance of the proposed method against other well-performing, global registration methods [15, 16, 17].
@@
easy group FGR [15] 0.98 0.93 59.8 22.23
Teaser++ [16] 0.99 0.93 60.4 22.74
Quatro [17] 1.05 1.04 49.5 23.37
@@
hard group FGR [15] 1.15 1.11 24.8 15.17
Teaser++ [16] 1.17 1.10 25.7 14.79
Quatro [17] 1.23 1.21 19.2 16.67
@@
As indicated in Table II, it is evident that the best current global registration methods struggle with the time complexity required for real-time processing when facing large-scale point clouds such as those in the DAIR-V2X dataset. Although these methods achieve near 50% success rate and obtain reasonably accurate initial calibration values on the dataset, they lack practical applicability. In contrast, the proposed V2I-Calib method significantly outperforms existing global point cloud registration algorithms in terms of runtime and achieves considerably higher accuracy in simple scenarios with lower complexity. In hard groups with more occlusions, the precision of our method is impacted due to the lower accuracy of detection results. Notably, the running times for the methods in the more challenging group are generally lower than those in the simpler group, which is attributable to the fewer shared targets in hard scenes, reducing the computational load.
@@ IV-C2 Ablation experiment
In this section, we aim to validate the effectiveness of the core affinity formulation module presented in Section III-A and the extrinsic parameter optimization module discussed in Section III-C through a series of ablation studies.
@@
Figure 4: Ablation Comparison of Affinity Strategies. Strategy 1 combines core and category affinities; Strategy 2 uses core affinity alone; Strategy 3 combines angle and category affinities; Strategy 4 integrates length with category affinities.
(a) Ideal Scenario
(b) Challenging Scenario
(c) Complex Scenario
(d) Complex Scenario
As indicated in Tables III and IV, the core affinity formulation module introduced in this study significantly enhances robustness, evidenced by a 20% increase in success rates when comparing V2I-Calib-v1 with V2I-Calib-v3. Furthermore, the inclusion of the extrinsic parameter optimization module, as evaluated through a comparison of V2I-Calib-v1 and V2I-Calib-v2, substantially improves calibration precision and success rates.
The superiority of the proposed affinity metrics was validated against category, angle, and length affinities, detailed in [8]. The results, depicted in Figure 4, show that Strategy 1 (core affinity combined with category affinity) not only achieves higher success rates and accuracy but also reduces computational time significantly compared to Strategy 2 (core affinity alone), especially in challenging scenarios.
Regarding the Strategy 3 (angle affinity with category affinity) and Strategy 4 (length affinity with category affinity) extended from [8], these strategies were originally designed for scenarios assisted by positional information. However, experimental results show that they are somewhat applicable in scenarios without initial extrinsic parameter values, although their performance significantly decreases in challenging settings. Upon further analysis of the scenarios where these strategies failed, we observed that most of these scenarios involved extrinsic parameters with relatively large absolute values, aligning with the limitations inherent in their initial design rationale.
@@ V CONCLUSIONS
This study introduces V2I-Calib, a calibration algorithm tailored for vehicle-infrastructure cooperative LiDAR systems. Unlike existing techniques, it does not require initial external parameters, enhancing stability in practical applications. Additionally, V2I-Calib satisfies real-time performance criteria, setting it apart from currently effective global point cloud registration methods.
Future research will focus on integrating temporal synchronization tasks to enhance practicality, alongside incorporating 2D bounding box data to improve multi-sensor fusion calibration. As the technology evolves, the V2I-Calib framework is expected to adapt to broader applications, enhancing the safety and reliability of autonomous driving systems.